OpenStudy (anonymous):

Solve the inequality 2y - 2 <14

OpenStudy (anonymous):

Simplifying 2y + -2 = 14 Reorder the terms: -2 + 2y = 14 Solving -2 + 2y = 14 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '2' to each side of the equation. -2 + 2 + 2y = 14 + 2 Combine like terms: -2 + 2 = 0 0 + 2y = 14 + 2 2y = 14 + 2 Combine like terms: 14 + 2 = 16 2y = 16 Divide each side by '2'. y = 8 Simplifying y=8

OpenStudy (anonymous):

Um…Thats not a choice!